OpenStudy (anonymous):

Using Mathematica, Find F's tangent line at a fixed point ( I already defined f(x) as (2x^2+32x)/(-x^3)... i just cannot for the life of me figure out the correct format

OpenStudy (anonymous):

f[x_]:=(function) f'[x]=(derivative) f'[x1]=(slope of the function at that point] Tangent line eqn= y-y1=m(x-x1) m=f'[x1] point (y1,x1)